# Diffusion Models for Tabular Data Imputation and Synthetic Data Generation

### v1

@@ 6.1.1. Synthetic data generation
The results shown in Table 2 indicate that MTabGen II outperforms existing state-of-the-art methods like TabDDPM, Tabsyn, CoDi, TVAE, or CTGAN in the synthetic data generation task, with the exception of the specialized MTabGen I which achieves better performance. The key outcomes of our experiments are as follows: (1) TVAEproduces better results than CTGAN. (2) Approaches based on Diffusion Models outperform TVAE and CTGAN in mean. (3) Our two proposed models systematically enhance the performance of TabDDPM, Tabsyn and CoDi, obtaining the best results across all datasets included in our benchmark. (4) Our model tends to outperform the baselines in datasets with a large number of features, such as the Gas Concentrations and Otto Group datasets, although this trend is less consistent with Tabsyn. Additionally, it is worth noting that the ML efficiency results align with those reported in the Tabsyn paper, where Tabsyn generally surpasses TabDDPM and CoDi, with TVAE also showing stronger performance than CoDi.
The results presented in Table 2 are obtained after applying an optimisation step for each generative model, using the Optuna library over $100$ trials, and evaluating performance with the cross-validated ML efficiency metric defined in Section 5.3 as the objective. The specific hyperparameters search space for each model is shown in Table 8, appendix A.
@@ 6.1.2. Ablation study
  • Enhanced learning of inter-feature interactions within condition (encoder) and masked features (decoder). Transformer layers allow for better learning of inter-feature interactions compared to MLPs. This is primarily due to the attention mechanism, where the new embedding of a feature $x_{i}$ is computed by a linear combination of the embeddings of all features $\{x_{j}\}$. The weight of feature $x_{j}$ depends on the current values of $x_{i}$ and $x_{j}$. This mechanism is more flexible than in MLP case, where the contribution of feature $x_{j}$ to a neuron in the next layer is fixed and does not depend on its current value.
@@ 6.1.3. Missing data imputation
Following the example of [60] and [65], the evaluation of imputed data quality is carried out in the Missing Completely at Random (MCAR) setup, under the assumption that the probability for missing data is uniformly distributed across observations and columns. To this end, we initially divide the dataset into two subsets: $70\%$ “imputation training” set and $30\%$ hold-out. Then, we construct three different versions of the $70\%$ “imputation training” split, each with increasing percentage of missing data: $10\%$, $25\%$, and $40\%$, while keeping the hold-out split intact. To evaluate the quality of imputed data in terms of ML efficiency, we apply the steps described in Alg. 1.
@@
2:     for each percentage of missing values in $\left\{10\%,25\%,40\%\right\}$ do
3:         Employ the generative model to fill in the missing values in the $70\%$ training subset.
4:         Train the ML discriminative model (e.g., XGBoost, CatBoost, LightGBM and MLP) on the imputed data.
5:         Evaluate the ML discriminative model on the hold-out $30\%$ subset.
6:     end for
7:end for
Our findings, detailed in Table 4, show that MTabGen consistently outperforms the baseline models considered in this step: missForest [48], GAIN [60], HyperImpute [23], Miracle [31], and TabCSDI [65]. Across all levels of missing data, MTabGen outperforms on average the baselines, and introduces a performance advantage that increases proportionally to the volume of missing data.
Table 4. Missing imputation analysis in terms of Machine Learning efficiency. Classification tasks use F1-score, and regression tasks use MSE, indicated by up/down arrows for maximization/minimization of the metric. Cross-validation mean and standard deviation are shown for each dataset-model pair. Best and second-best results are highlighted in bold and underline, respectively.
@@ 6.2. Statistical Similarity
Table 5-a shows the average Wasserstein Distance between synthetic and real numerical data distributions. Specifically, the Wasserstein distance is calculated for each numerical column between the real data and the synthetic data generated by Tabsys, TabDDPM, and MTabGen. For each generative model, the final dataset results are the average of these distances across all numerical columns in the dataset. MTabGen consistently performs better than Tabsys and TabDDPM. This advantage is more pronounced in datasets where there is a larger difference in ML efficiency, such as HELOC, California Housing, House Sales or Forest Cover Type dataset. In contrast, in datasets like Insurance, where all models have similar ML efficiency, the Wasserstein distances are also comparable. A qualitative analysis of the results for some of the columns can be found in the four plots in the two top rows of Fig. 4. These plots compare the distributions of real data with those of the synthetic data generated by the different models.
Similarly, Table 5-b shows the average Jensen-Shannon distance between synthetic and real categorical data distributions. MTabGen and Tabsyn outperform TabDDPM, consistent with the findings of [62], where Tabsyn achieved better results than TabDDPM on categorical data. Also the Jensen-Shannon distance appears to be strongly related to ML efficiency. For example, in datasets like HELOC, House Sales or Forest Cover Type, where ML efficiency of MTabGen is significantly better than the other baselines, it also has a lower Jensen-Shannon distance. In the Insurance dataset, where all baselines have similar ML efficiency, the Jensen-Shannon distances are also similar. Notably, while MTabGen and Tabsyn show similar overall results, a qualitative analysis of per-column results reveals that MTabGen tends to excel when the number of classes in the categorical variable increases, as shown in the bottom row plots of Fig. 4.
Table 5-c shows the average L2 distance between the two correlation matrices computed on real and synthetic data. Fig. 5, instead, shows the L2 distance details across all the datasets in the benchmark. In Fig. 5 more intense green color means higher difference between the real and synthetic correlation values. Even more than in the previous statistical measures, there exist an association between ML efficiency and L2 distance between correlation matrices. In datasets like HELOC, California Housing, House Sales or Forest Cover Type where the ML efficiency of MTabGen is notably better than the one of Tabsyn and TabDDPM the corresponding heatmap of MTabGen are more lighter, e.g. they shown a smaller error in the correlation estimation. Nevertheless in datasets like Cardio or Insurance where all the models perform similarly, the heatmaps do not shown resignable differences.
@@ 6.3. Privacy risk
In this section, we delve deeper into the privacy risk associated with synthetic data. As previously mentioned, DCR is defined by the Euclidean distance between any synthetic record and its closest real neighbor. Ideally, a higher DCR indicates a lower privacy risk. However, out-of-distribution data (random noise) can also result in high DCR. Therefore, to maintain ecological validity, DCR should be evaluated alongside the ML efficiency metric. For this reason, our privacy risk evaluation includes only Tabsyn, TabDDPM, and MTabGen. These models have superior ML efficiency and are more likely to pose a privacy risk (lower DCR) because they closely mimic the original data
Figures 6 and 7 provide a qualitative comparison of the DCR distribution for MTabGen versus Tabsyn and TabDDPM when the relative improvement in ML efficiency is less than 5%. This comparison confirms that there are no significant changes or patterns, indicating that the improvement in ML efficiency does not lead to a noticeable increase in privacy risk.
@@
Figure 6. Distance to Closest Record (DCR) for Tabsyn and MTabGen
Figure 7. Distance to Closest Record (DCR) for TabDDPM and MTabGen
@@ 7. Conclusion
In this paper, we introduced MTabGen, a diffusion model enhanced with a conditioning attention mechanism, a transformer-based encoder-decoder architecture, and dynamic masking. MTabGen is specifically designed for applications involving mixed-type tabular data. The transformer encoder-decoder acts as the denoising network, enabling the conditioning attention mechanism while effectively capturing and representing complex interactions and dependencies within the data. The dynamic masking feature allows MTabGen to handle both synthetic data generation and missing data imputation tasks within a unified framework efficiently. We proposed to train the diffusion model to regenerate masked data, enabling applications ranging from data imputation to unconditioned or conditioned synthetic data generation. This versatility makes MTabGen suitable for generating synthetic data to overcome privacy regulations, augment existing datasets, or mitigate class imbalances. We evaluated MTabGen against established baselines across several public datasets with a diverse range of features. Our model demonstrated better performance in terms of ML efficiency and statistical accuracy, while maintaining privacy risks comparable to those of the baselines, particularly showing increased performance in datasets with a large number of features.

### v2

@@ 6.1.1. Synthetic data generation
The results shown in Table 2 indicate that MTabGen II demonstrates competitive performance compared to existing state-of-the-art methods like TabDDPM, Tabsyn, CoDi, TVAE, or CTGAN in the synthetic data generation task, while showing moderate but statistically significant7 improvements in most of the datasets. However, the specialized MTabGen I achieves the best performance across the evaluated tasks. The key outcomes of our experiments are as follows: (1) TVAEproduces better results than CTGAN. (2) Approaches based on Diffusion Models outperform TVAE and CTGAN on average. (3) Our two proposed models show overall improved performance over TabDDPM, Tabsyn and CoDi, with statistically significant improvements in specific datasets777 Following the recommendation of [46], we applied the Wilcoxon signed-rank test to compare, for each dataset, the results obtained by the best-performing MTabGen model against the best-performing baseline. In 6 out of 10 cases-specifically, HELOC, California Housing, House Sales, Adult Income, Otto, and Forest Cover Type, MTabGen demonstrated a statistically significant improvement over the strongest baseline, with a p-value $<0.01$.. In the remaining cases, the results were on par with the baselines, without a statistically significant difference. (4) Our model tends to outperform the baselines in datasets with a large number of features, such as the Gas Concentrations and Otto Group datasets, although this trend is less consistent with Tabsyn. Additionally, it is worth noting that the ML efficiency results align with those reported in the Tabsyn paper, where Tabsyn generally surpasses TabDDPM and CoDi, with TVAE also showing stronger performance than CoDi.
The results presented in Table 2 are obtained after applying Bayesian optimisation for each generative model, using the Optuna library over $100$ trials, and evaluating performance with the cross-validated ML efficiency metric defined in Section 5.3 as the objective. A similar hyperparameter optimization procedure was applied to all baseline models, considering the parameters specific to each method, enabling fair comparison with MTabGen. The specific hyperparameter search space for each model is shown in Table 10, Appendix A.
@@ 6.1.2. Ablation study
  • Enhanced learning of inter-feature interactions within condition (encoder) and masked features (decoder). Transformer layers allow for better learning of inter-feature interactions compared to MLPs. This is primarily due to the attention mechanism, where the new embedding of a feature $x_{i}$ is computed by a linear combination of the embeddings of all features $\left\{x_{j}\right\}$. The weight of feature $x_{j}$ depends on the current values of $x_{i}$ and $x_{j}$. This mechanism is more flexible than in MLP case, where the contribution of feature $x_{j}$ to a neuron in the next layer is fixed and does not depend on its current value.
@@ 6.1.3. Missing data imputation
[…]
3:         for each percentage of missing values in $\left\{10\%,25\%,40\%\right\}$ do
4:              Train the generative model on $30\%$ “imputation training” subset.
5:              Impute the missing values in the $30\%$ “imputation testing/discriminative training” subset.
6:              Train the ML discriminative model (e.g., XGBoost, CatBoost, LightGBM and MLP) on the imputed data.
7:              Evaluate the ML discriminative model on the hold-out $30\%$ subset.
8:         end for
9:     end for
10:end for
The evaluation of imputed data quality considers three possible scenarios for missing data: Missing Completely at Random (MCAR), Missing at Random (MAR), and Missing Not at Random (MNAR). For each scenario, we first divide the dataset into three subsets: $40\%$ for “imputation training”, $30\%$ for “imputation testing/discriminative training” and $30\%$ as a hold-out set. Next, we generate three versions of the imputation training and imputation test/discriminative train splits, each containing an increasing proportion of missing data ($10\%$, $25\%$, and $40\%$), while keeping the hold-out set intact. To assess the quality of the imputed data in terms of machine learning efficiency, we follow the steps outlined in Alg. 3. For the MAR and MNAR scenarios, the missing data are generated using the code used by [25] and [39].
Table 4. Missing imputation analysis in terms of Machine Learning efficiency when the missing information is MCAR. Classification tasks use F1-score, and regression tasks use MSE, indicated by up/down arrows for maximization/minimization of the metric. Cross-validation mean and standard deviation are shown for each dataset-model pair. Best and second-best results are highlighted in bold and underline, respectively.
@@
Table 5. Missing imputation analysis in terms of Machine Learning efficiency when the missing information is MAR. Classification tasks use F1-score, and regression tasks use MSE, indicated by up/down arrows for maximization/minimization of the metric. Cross-validation mean and standard deviation are shown for each dataset-model pair. Best and second-best results are highlighted in bold and underline, respectively.
[…]
$10\%$ $83.13\pm 0.02$ $82.39\pm 0.04$ $83.34\pm 0.02$ $83.30\pm 0.03$ $83.25\pm 0.03$ $\mathbf{83.78\pm 0.02}$
HELOC $\uparrow$ $25\%$ $82.75\pm 0.03$ $81.88\pm 0.02$ $83.15\pm 0.03$ $82.81\pm 0.02$ $83.18\pm 0.01$ $\mathbf{83.53\pm 0.04}$
$40\%$ $82.42\pm 0.02$ $81.64\pm 0.04$ $82.86\pm 0.04$ $82.52\pm 0.02$ $82.59\pm 0.05$ $\mathbf{83.40\pm 0.04}$
$10\%$ $84.46\pm 0.04$ $84.55\pm 0.03$ $84.42\pm 0.02$ $84.58\pm 0.04$ $84.44\pm 0.03$ $\mathbf{84.76\pm 0.03}$
Churn $\uparrow$ $25\%$ $84.05\pm 0.03$ $83.39\pm 0.02$ $83.79\pm 0.02$ $84.00\pm 0.02$ $83.58\pm 0.02$ $\mathbf{84.50\pm 0.02}$
$40\%$ $82.91\pm 0.05$ $83.11\pm 0.01$ $83.00\pm 0.03$ $82.40\pm 0.04$ $83.31\pm 0.03$ $\mathbf{84.17\pm 0.04}$
$10\%$ $0.182\pm 0.003$ $0.183\pm 0.004$ $0.173\pm 0.003$ $0.174\pm 0.003$ $0.175\pm 0.004$ $\mathbf{0.171\pm 0.003}$
Cal. Hous. $\downarrow$ $25\%$ $0.230\pm 0.002$ $0.251\pm 0.004$ $0.191\pm 0.002$ $0.197\pm 0.004$ $0.190\pm 0.005$ $\mathbf{0.183\pm 0.002}$
$40\%$ $0.272\pm 0.002$ $0.308\pm 0.005$ $0.222\pm 0.003$ $0.274\pm 0.002$ $0.264\pm 0.004$ $\mathbf{0.211\pm 0.003}$
$10\%$ $0.110\pm 0.004$ $0.112\pm 0.002$ $0.114\pm 0.004$ $0.106\pm 0.002$ $0.111\pm 0.004$ $\mathbf{0.105\pm 0.002}$
House Sales $\downarrow$ $25\%$ $0.120\pm 0.003$ $0.134\pm 0.002$ $0.122\pm 0.004$ $0.130\pm 0.002$ $0.121\pm 0.003$ $\mathbf{0.116\pm 0.003}$
$40\%$ $0.158\pm 0.003$ $0.149\pm 0.001$ $0.160\pm 0.003$ $0.134\pm 0.003$ $0.132\pm 0.003$ $\mathbf{0.120\pm 0.002}$
Table 6. Missing imputation analysis in terms of Machine Learning efficiency when the missing information is MNAR. Classification tasks use F1-score, and regression tasks use MSE, indicated by up/down arrows for maximization/minimization of the metric. Cross-validation mean and standard deviation are shown for each dataset-model pair. Best and second-best results are highlighted in bold and underline, respectively.
[…]
$10\%$ $83.09\pm 0.03$ $82.35\pm 0.03$ $83.31\pm 0.04$ $83.25\pm 0.04$ $83.23\pm 0.04$ $\mathbf{83.75\pm 0.04}$
HELOC $\uparrow$ $25\%$ $82.59\pm 0.03$ $81.78\pm 0.02$ $83.04\pm 0.05$ $82.69\pm 0.05$ $83.07\pm 0.05$ $\mathbf{83.47\pm 0.03}$
$40\%$ $82.12\pm 0.04$ $81.43\pm 0.03$ $82.67\pm 0.03$ $82.35\pm 0.03$ $82.31\pm 0.04$ $\mathbf{83.31\pm 0.04}$
$10\%$ $84.42\pm 0.04$ $84.51\pm 0.05$ $84.36\pm 0.04$ $84.49\pm 0.03$ $84.38\pm 0.04$ $\mathbf{84.72\pm 0.04}$
Churn $\uparrow$ $25\%$ $83.95\pm 0.04$ $83.27\pm 0.04$ $83.66\pm 0.05$ $83.86\pm 0.04$ $83.72\pm 0.03$ $\mathbf{84.42\pm 0.03}$
$40\%$ $82.72\pm 0.03$ $82.93\pm 0.05$ $82.80\pm 0.05$ $82.23\pm 0.03$ $83.09\pm 0.04$ $\mathbf{84.04\pm 0.05}$
$10\%$ $0.184\pm 0.003$ $0.185\pm 0.004$ $0.177\pm 0.004$ $0.178\pm 0.003$ $0.176\pm 0.003$ $\mathbf{0.173\pm 0.003}$
Cal. Hous. $\downarrow$ $25\%$ $0.234\pm 0.004$ $0.255\pm 0.003$ $0.200\pm 0.003$ $0.205\pm 0.002$ $0.194\pm 0.004$ $\mathbf{0.187\pm 0.003}$
$40\%$ $0.285\pm 0.003$ $0.318\pm 0.002$ $0.239\pm 0.004$ $0.289\pm 0.003$ $0.279\pm 0.002$ $\mathbf{0.220\pm 0.002}$
$10\%$ $0.111\pm 0.002$ $0.116\pm 0.004$ $0.117\pm 0.004$ $0.112\pm 0.002$ $0.115\pm 0.003$ $\mathbf{0.108\pm 0.003}$
House Sales $\downarrow$ $25\%$ $0.127\pm 0.001$ $0.140\pm 0.002$ $0.129\pm 0.003$ $0.137\pm 0.004$ $0.126\pm 0.002$ $\mathbf{0.120\pm 0.002}$
$40\%$ $0.169\pm 0.002$ $0.160\pm 0.003$ $0.171\pm 0.003$ $0.143\pm 0.003$ $0.140\pm 0.004$ $\mathbf{0.126\pm 0.003}$
The results of the experiments for the MCAR, MAR, and MNAR scenarios are presented in Tables 4, 5, 6, respectively. As the assumption about missingness becomes more complex (MCAR $\rightarrow$ MAR $\rightarrow$ MNAR), all the models included in the comparison exhibit a very slight degradation in performance. However, the findings remain consistent across all missing data scenarios.
Our experiments indicate that MTabGen tends to outperform the baseline models considered in this evaluation: missForest [53], GAIN [65], HyperImpute [25], Miracle [33], and TabCSDI [70]. Across all levels of missing data, MTabGen generally achieves better performance on average, with its advantage becoming more pronounced as the proportion of missing data increases.
@@ 6.2. Statistical Similarity
Table 7-a shows the average Wasserstein Distance between synthetic and real numerical data distributions. Specifically, the Wasserstein distance is calculated for each numerical column between the real data and the synthetic data generated by Tabsys, TabDDPM, and MTabGen. For each generative model, the final dataset results are the average of these distances across all numerical columns in the dataset. MTabGen consistently performs better than Tabsys and TabDDPM. This advantage is more pronounced in datasets where there is a larger difference in ML efficiency, such as HELOC, California Housing, House Sales or Forest Cover Type dataset. In contrast, in datasets like Insurance, where all models have similar ML efficiency, the Wasserstein distances are also comparable. A qualitative analysis of the results for some of the columns can be found in the four plots in the first two rows plots of Fig. 4. These plots compare the distributions of real data with those of the synthetic data generated by the different models.
Similarly, Table 7-b shows the average Jensen-Shannon distance between synthetic and real categorical data distributions. MTabGen and Tabsyn outperform TabDDPM, consistent with the findings of [67], where Tabsyn achieved better results than TabDDPM on categorical data. Also the Jensen-Shannon distance appears to be strongly related to ML efficiency. For example, in datasets like HELOC, House Sales or Forest Cover Type, where ML efficiency of MTabGen is significantly better than the other baselines, it also has a lower Jensen-Shannon distance. In the Insurance dataset, where all baselines have similar ML efficiency, the Jensen-Shannon distances are also similar. Notably, while MTabGen and Tabsyn show similar overall results, a qualitative analysis of per-column results reveals that MTabGen tends to excel when the number of classes in the categorical variable increases, as shown in the last two rows plots of Fig. 4.
@@
Table 7-c shows the average L2 distance between the two correlation matrices computed on real and synthetic data. Fig. 5, instead, shows the L2 distance details across all the datasets in the benchmark. In Fig. 5 more intense green color means higher difference between the real and synthetic correlation values. Even more than in the previous statistical measures, there exist an association between ML efficiency and L2 distance between correlation matrices. In datasets like HELOC, California Housing, House Sales or Forest Cover Type where the ML efficiency of MTabGen is notably better than the one of Tabsyn and TabDDPM the corresponding heatmap of MTabGen are more lighter, e.g. they shown a smaller error in the correlation estimation. Nevertheless in datasets like Cardio or Insurance where all the models perform similarly, the heatmaps do not show resignable differences.
@@ 6.3. Privacy risk
In this section, we delve deeper into the privacy risk associated with synthetic data. As previously mentioned, DCR is defined by the Euclidean distance between any synthetic record and its closest real neighbor. Ideally, a higher DCR indicates a lower privacy risk. However, out-of-distribution data (random noise) can also result in high DCR. Therefore, to maintain ecological validity, DCR should be evaluated alongside the ML efficiency metric. For this reason, our privacy risk evaluation includes only Tabsyn, TabDDPM, and MTabGen. These models have superior ML efficiency and are more likely to pose a privacy risk (lower DCR) because they closely mimic the original data.
@@
Figure 6. Distance to Closest Record (DCR) for Tabsyn and MTabGen.
Figure 7. Distance to Closest Record (DCR) for TabDDPM and MTabGen.
Figures 6 and 7 provide a qualitative comparison of the DCR distribution for MTabGen versus Tabsyn and TabDDPM when the relative improvement in ML efficiency is less than 5%. Specifically, for each synthetic data point, we compute the distance to its closest record in the training dataset. The plots then depict the distribution of these distances. This comparison confirms that there are no significant changes or patterns, indicating that the improvement in ML efficiency does not lead to a noticeable increase in privacy risk.
@@ 7. Conclusion
In this paper, we introduced MTabGen, a diffusion model enhanced with a conditioning attention mechanism, a transformer-based encoder-decoder architecture, and dynamic masking. MTabGen is specifically designed for applications involving mixed-type tabular data. The transformer encoder-decoder acts as the denoising network, enabling the conditioning attention mechanism while effectively capturing and representing complex interactions and dependencies within the data. The dynamic masking feature allows MTabGen to handle both synthetic data generation and missing data imputation tasks within a unified framework efficiently. We proposed to train the diffusion model to regenerate masked data, enabling applications ranging from data imputation to unconditioned or conditioned synthetic data generation. This versatility makes MTabGen suitable for generating synthetic data to overcome privacy regulations, augment existing datasets, or mitigate class imbalances. We evaluated MTabGen against established baselines across several public datasets with a diverse range of features. Our model demonstrated better overall performance in terms of ML efficiency and statistical accuracy, while maintaining privacy risks comparable to those of the baselines, particularly showing increased performance in datasets with a large number of features.
@@ Appendix A Hyperparameter tuning
Appendix B Addition Experimental Results
In this Appendix, we compare the training and sampling times, as well as the number of trainable parameters, of MTabGen against other tabular generative models, using the Churn, California Housing and Gas Concentrations datasets as examples. The training and sampling times and the number of trainable parameters are highly influenced by both the model hyperparameters and the dataset characteristics (size and number of features).
Regarding model hyperparameters, each model has been optimized for these datasets as outlined in Section 6.1.1. This process involves using Bayesian optimization and Optuna to identify the best hyperparameters for each model, enhancing the machine learning efficiency described in Section 5.3.1. The specific hyperparameter search space for each model is provided in Table 10 in Appendix A, and Table 11 presents the training and sampling times, along with the number of trainable parameters, for the optimized models on the Churn, California Housing and Gas Concentrations datasets.
In terms of dataset characteristics, Churn, California Housing and Gas Concentrations differ in size ($10000$ vs. $20640$ vs. $13910$ samples) and feature count ($10$ vs. $8$ vs. $129$), impacting absolute values such as training and sampling times, as well as the number of trainable parameters. Despite these differences, all the datasets yield consistent results.
Table 11. Taining/Sampling time and number of trainable parameters comparison for Churn, California Housing and Gas Concentrations datasets.
Churn California Housing Gas Concentrations
Training Sampling Parameters Training Sampling Parameters Training Sampling Parameters
TVAE 5m 3s 0.86s 190122 8m 10s 1.96s 180132 14m 47s 4.35s 325108
CTGAN 14m 25s 1.68s 280121 18m 36s 3.01s 265132 29m 58s 5.21s 456587
CoDi 2h 26m 5s 6.64s 615212 2h 54m 10s 10.54s 567987 3h 58m 15s 16.38s 953578
Tabsyn 31m 34s 6.02s 415122 38m 42s 9.16s 409876 1h 5m 31s 14.43s 762463
TabDDPM 14m 36s 12.75s 221363 18m 12s 18.21s 209815 29m 36s 28.48s 420589
MTabGen I 14m 42s 15.12s 332049 18m 10s 21.10s 298844 31m 48s 30.45s 597688
MTabGen II 14m 53s 15.30s 332049 18m 42s 22.95s 298844 32m 24s 35.12s 597688
From the results in Table 2, the baseline model that most closely matches the performance of MTabGen I and MTabGen II in terms of machine learning efficiency is Tabsyn. Tabsyn is a latent diffusion model where a VAE first projects tabular data into a dense, homogeneous continuous space, followed by a diffusion process. Compared to our model, Tabsyn has a greater number of trainable parameters and longer training times. However, in terms of sampling, Tabsyn outperforms our model. This advantage largely arises from the findings of [28], which suggest that the sampling process can be accelerated by reducing the number of backward diffusion steps through an effective choice of schedule and scale functions (see [28] for details). The schedule function determines the desired noise level at each diffusion time step, while the scale function dictates how data scales over time. Their optimal selection aligns with that proposed by DDIM [51]. As highlighted by [28], this choice is independent of how the denoising model (e.g., the neural network) has been trained. Therefore, we plan to investigate the integration of our masking training mechanism with the sampling approach proposed by [28] in future work to improve the model’s sampling speed.
A final comment addresses the impact of dynamic conditioning, specifically the differences between MTabGen I and MTabGen II. Table 11 demonstrates that the conditioning mechanism has a limited effect on sampling and training time, while the number of trainable parameters remains unchanged with or without conditioning.
This outcome arises because the conditioning mask only influences the input size of the encoder-decoder transformer, specifically affecting the sequence length of the transformer encoder input (as illustrated in Figure 2 and discussed in Section 4.3). Notably, the number of trainable parameters in a transformer encoder does not depend on input length; therefore, MTabGen I and MTabGen II both retain the same parameter count. Although input sequence length can impact computational time, this effect is minimal in our case. Even with the largest dataset, consisting of 129 features, the encoder input remains a relatively short sequence of 129 tokens. Consequently, the impact on both sampling and training time is minimal.